### v1

@@ Abstract
Purpose: Ultrasound-based bone surface segmentation is crucial in computer-assisted orthopedic surgery. However, ultrasound images have limitations, including a low signal-to-noise ratio, acoustic shadowing, and speckle noise, which make interpretation difficult. Existing deep learning models for bone segmentation rely primarily on costly manual labeling by experts, limiting dataset size and model generalizability. Additionally, the complexity of ultrasound physics and acoustic shadow makes the images difficult for humans to interpret, leading to incomplete labels in low-intensity and anechoic regions and limiting model performance. To advance the state of the art in ultrasound bone segmentation and establish effective model benchmarks, larger and higher-quality datasets are needed.
Methods: We propose a methodology for collecting ex-vivo ultrasound datasets with automatically generated bone labels, including anechoic regions. The proposed labels are derived by accurately superimposing tracked bone Computed Tomography (CT) models onto the tracked ultrasound images. These initial labels are refined to account for ultrasound physics. A clinical evaluation is conducted by an expert physician specialized on orthopedic sonography to assess the quality of the generated bone labels. A neural network for bone segmentation is trained on the collected dataset and its predictions are compared to expert manual labels, evaluating accuracy, completeness, and F1-score.
Results: We collected the largest known dataset of 100k ultrasound images of human lower limbs with bone labels, called UltraBones100k. A Wilcoxon signed-rank test with Bonferroni correction confirmed that the bone alignment after our optimization pipeline significantly improved the quality of bone labeling ($p<0.001$). The model trained on UltraBones100k consistently outperforms manual labeling in all metrics, particularly in low-intensity regions (320% improvement in completeness at a distance threshold of 0.5 mm).
@@
Bone segmentation, Bone surface segmentation, Bone annotation, Ultrasound image analysis, Computer-assisted orthopedic surgery
@@ 1 Introduction
Ultrasound is a well-established imaging technology widely used in the clinical diagnosis of musculoskeletal pathologies, including bone and soft tissue injuries [1]. It also facilitates dynamic assessments of joint and tendon movements, stability[2], and the detection of fracture healing, structural abnormalities, infections, ligament injuries, nerve compression, and mechanical impingement [3]. As evidenced by a substantial body of literature [4, 5, 6, 7, 8], the segmentation and 3D reconstruction of ultrasound are gaining increasing significance as a powerful modality in interventional Computer-Assisted Orthopedic Surgery (CAOS) systems. Compared to ionizing imaging modalities like fluoroscopy, ultrasound offers advantages in improving guidance and navigation during interventional and surgical procedures[4], including trauma treatment[6], needle placement, and robotic surgery [5].
@@
Accurate 3D modeling of bone geometry requires precise identification and segmentation of bone anatomies, which relies heavily on high-quality labels produced by expert professionals. Due to the highest acoustic impedance compared to surrounding soft tissues, bones typically appear as high-intensity regions when the ultrasound probe is perpendicular to the bone surface. This phenomenon holds true across various bones and individuals, facilitating robust computer-based segmentation of bones in ultrasound images. However, when the probe is positioned at a larger oblique angle to the bone surface, bones can appear as low-intensity regions as most portion of acoustic energy is reflected into surrounding soft tissues instead of back to the probe. These low-intensity regions can only be inferred by surgeons through anatomical knowledge but cannot be accurately labeled. The same applies to fully anechoic regions formed by the complete absorption of the ultrasound signal in the overlying bone layers.
@@
To address these challenges, various medical imaging methods based on handcrafted features have been developed to enhance the interpretation of ultrasound images [9, 10]. Recently, deep learning-based bone segmentation methods have shown promising improvements in robustness, accuracy, and processing speed [11, 7, 8, 12, 13]. However, the performance of such models relies on the quality of the labeled training dataset, which cannot be produced on a large scale if relying solely on human experts. The resulting datasets are often small (ranging from dozens to thousands labeled slices), limiting the performance and generalization of the trained models. Additionally, most existing methods train and validate their models on non-public datasets and use non-standardized evaluation metrics, making it difficult to compare and benchmark the performance of different models across studies [14, 15]. The absence of large-scale, accurately labeled public datasets has been a critical obstacle for the community in developing more advanced models.
@@
In this work, we propose a methodology for collecting ultrasound datasets with automatically generated bone labels. Instead of relying on manual labeling, our approach utilizes tracked CT bone model to tackle the problem of identifying obscured, low-intensity bone regions, resulting in more precise and complete labels. With this methodology, we have collected a dataset of 100k annotated ultrasound images with bone labels from 14 human cadaveric lower legs, called UltraBones100k. We hypothesize that the information in UltraBones100k is sufficient for deep learning models to segment both high intensity and low intensity bone regions. A deep learning model for bone segmentation is trained on UltraBones100k, which significantly outperforms manual segmentation performed by an expert surgeon.
The dataset, as well as the trained model, is available for download here [^1] Link will be available upon acceptance..
@@ 2 Related Work
Ultrasound Bone Segmentation. Numerous intensity-based bone segmentation methods have been developed [14, 15]. However, intensity-based methods face challenges in terms of robustness and accuracy. To mitigate these issues, frequency-based methods have been introduced, including phase symmetry and structured phase symmetry techniques [10, 9], which aim to enhance bone boundaries. With the advent of deep learning, neural network models such as CNNs, U-nets, and GANs have been employed for bone segmentation in ultrasound images, targeting various anatomical regions such as the spine [11, 7], pelvis [8], femur [12, 8, 13], tibia [8, 13, 12], fibula [12], and radius [13, 12]. These deep learning methods have demonstrated superior performance over traditional approaches in terms of accuracy, robustness, and runtime. For example, average bone surface distance errors have been reported to range from 0.2 mm to 1.7 mm [13, 16, 7]. Based on accurate 2D bone segmentation results, the reconstructed bone surfaces achieve sub-millimetric one-sided CD [17]. However, these studies rely on small non-public datasets with manually labeled bone structures and fail to use effective separate validation and test sets, limiting the generalizability of trained models. The absence of standardized evaluation metrics, with over 18 different metrics in use [14], further hinders effective model benchmarking and comparison.
Automatic labeling. Optical tracking markers have been applied in the medical imaging community for collecting datasets with automatically generated labels, such as tracked spine CT models [18] and surgical instruments [19]. So far, one previous work has attempted to automatically generate bone labels in ultrasound images using a tracked bone CT model [20]. However, their approach led to significant alignment errors due to error accumulation, ultimately requiring a fallback to manual labeling. Image-based multi-modal registration methods, such as CT-US registration in Imfusion Suite  [^2] https://www.imfusion.com/, can also be used to automatically align CT-US data. However, as we will demonstrate, these methods do not provide sufficient accuracy for automatic labeling.
@@ 3.1 Dataset Collection and Automatic Bone labeling
Acquisition Setup. Our dataset acquisition setup using human cadaveric specimens is illustrated in Figure 1.333Ethics approval has been granted by the Zurich cantonal ethics committee under BASEC Nr. 2023-01681. Prior to data acquisition, K-wires with a diameter of 2.5 mm and a length of 150 mm (DePuy Synthes, USA), along with custom-made 3D-printed marker mounts equipped with infrared-sensitive fiducial markers, were mounted to each specimen. The K-wires were used to stabilize the bones of interest, including the lower leg and foot, to prevent bone movement during scanning. Optical markers were tracked by an optical tracking system (FusionTrack 500, Atracsys, Switzerland). A CT scan was performed on each specimen with an image resolution of 512 × 512 pixels, an in-plane resolution of 0.839 mm × 0.839 mm, and a slice thickness of 0.6 mm (NAEOTOM Alpha, Siemens, Germany). Bone segmentation was generated using the region growing and thresholding functionalities in Mimics (Materialise, Leuven, Belgium). For data acquisition, each specimen was scanned nine times using a tracked ultrasound device (AixplorerUltimate, SuperSonic Imagine, France) equipped with a linear probe that has an imaging frequency range of 4.0 to 15.0 MHz. The scanning procedure consisted of three forward and backward passes along the longitudinal axis of the tibia, three along the fibula, and three covering the foot (including the bones of the ankle joint such as the calcaneus and talus), aiming to maximize bone surface coverage. Ultrasound images, along with the tracking data of the ultrasound probe and the specimen, were recorded with corresponding timestamps.
@@
System Calibration. To determine the coordinates of the tracking fiducials on the specimen marker within the CT space, virtual spheres are created to match the size of the physical fiducials in MeshLab  [^4] https://www.meshlab.net/. These virtual spheres are initially manually aligned with the corresponding fiducials in the segmented CT model, followed by refinement using the point-to-point Iterative Closest Point algorithm [21]. The resulting registration transformation matrix allows for the accurate calculation of relative poses between the ultrasound probe and the bone CT model. However, converting each pixel of an ultrasound image into its corresponding point in the CT space requires accurate spatial calibration and accurate tracking of the ultrasound probe. In addition, temporal calibration is necessary to synchronize the ultrasound images with the tracking data. Following the convention in existing works [22, 23], we define the origin of the ultrasound image at the top-left corner. The pixels of the ultrasound image at timestamp $t$ are defined in homogeneous coordinates as $\mathcal{P}^{t}=\left\{\begin{bmatrix}u,v,0,1\end{bmatrix}^{T}\mid u\in[0,W),v\in[0,H)\right\}$, where $W$ and $H$ represent the image width and height, respectively. The corresponding point cloud in CT space, $\mathcal{U}^{t}=\left\{\begin{bmatrix}x_{j},y_{j},z_{j},1\end{bmatrix}^{T}\mid j\in[0,H\times W)\right\}$, is computed as
$\mathcal{U}^{t}=\mathsf{T}^{t+\Delta t}_{CT\leftarrow O}\cdot\mathsf{T}^{t+\Delta t}_{O\leftarrow US}\cdot\mathsf{T}_{US\leftarrow I}\cdot\mathsf{T}_{S}\cdot\mathcal{P}^{t}=\mathsf{T}^{t+\Delta t}_{CT\leftarrow US}\cdot\mathsf{T}_{US\leftarrow I}\cdot\mathsf{T}_{S}\cdot\mathcal{P}^{t}$ (1)
where
$\mathsf{T}_{US\leftarrow I}=\begin{bmatrix}r_{11}&r_{12}&r_{13}&d_{x}\\ r_{21}&r_{22}&r_{23}&d_{y}\\ r_{31}&r_{32}&r_{33}&d_{z}\\ 0&0&0&1\end{bmatrix},\quad\mathsf{T}_{S}=\begin{bmatrix}s_{x}&0&0&0\\ 0&s_{y}&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}.$ (2)
@@
Here, $\mathsf{T}_{S}$ transforms 2D pixels into 3D points in the ultrasound image coordinate system, where $s_{x}$ and $s_{y}$ are the pixel dimensions in mm, as specified by the device. The transformation matrix $\mathsf{T}_{US\leftarrow I}\in\text{SE}(3)$ (the Special Euclidean group in 3D) maps points from the image plane to the coordinate system of the ultrasound probe marker. Note that a transformation matrix $\mathsf{T}(\boldsymbol{\theta},\mathbf{d})$ can be defined using the Euler rotation angles $\boldsymbol{\theta}$ and the translation vector $\mathbf{d}$, which together specify the relationship between two coordinate systems. Similarly, $\mathsf{T}^{t+\Delta t}_{O\leftarrow US}\in\text{SE}(3)$ and $\mathsf{T}^{t+\Delta t}_{CT\leftarrow O}\in\text{SE}(3)$ represent transformations from the ultrasound marker to the optical tracking system and from the optical system to the CT space at timestamp $t+\Delta t$, respectively. The parameters of $\mathsf{T}^{t+\Delta t}_{O\leftarrow US}$ and $\mathsf{T}^{t+\Delta t}_{CT\leftarrow O}$ are provided by the optical tracking system. For simplicity, we denote $\mathsf{T}^{t+\Delta t}_{CT\leftarrow US}(\boldsymbol{\theta},\mathbf{d})=\mathsf{T}^{t+\Delta t}_{CT\leftarrow O}(\boldsymbol{\theta}_{1},\mathbf{d}_{1})\cdot\mathsf{T}^{t+\Delta t}_{O\leftarrow US}(\boldsymbol{\theta}_{2},\mathbf{d}_{2})$ as the direct transformation from the ultrasound marker to the CT space. The temporal offset $\Delta t$ compensates for lag between tracking and ultrasound signals [22]. $\mathsf{T}_{US\leftarrow I}$ is determined via spatial calibration [23]. Unlike previous studies that treated spatial and temporal calibration independently [23, 22], we performed both calibrations simultaneously using a Z-phantom [23]. The temporal offset $\Delta t$ was sampled over the range $[-300,0)$ ms with a step size of 0.1 ms. For each $\Delta t$, $\mathsf{T}_{US\leftarrow I}$ was computed, and the corresponding 3D localization recorded. The optimal parameters were identified by minimizing the 3D localization error. Our method achieved a 3D localization error of 1.07 mm, outperforming previously reported errors ($\approx 1.5$ mm) [23, 22]. With precise calibration results, 2D pixel coordinates can be accurately mapped to 3D points in the CT space. For simplicity, we denote the pixel corresponding to the point $\mathbf{u}_{i}^{t}\in\mathcal{U}^{t}$ as $\mathbf{p}_{i}^{t}\in\mathcal{P}^{t}$, and define $f(\mathbf{u}_{i}^{t})$ as the intensity value of $\mathbf{p}_{i}^{t}$ in the original ultrasound image.
@@
Bone Labeling. Points are sampled uniformly over the bone CT model to generate the CT point cloud $\mathcal{C}$. Both $\mathcal{C}$ and $\mathcal{U}^{t}$ (with original intensity values) are rendered in Figure 2. The CT-ultrasound bone surface intersection is defined as $\mathcal{I}^{t}=\mathcal{C}\cap\mathcal{U}^{t}=\left\{\mathbf{u}\in\mathcal{U}^{t}\mid\min_{\mathbf{c}\in\mathcal{C}}\lVert\mathbf{u}-\mathbf{c}\rVert\leq\gamma\right\}$ where $\gamma$ is the parameter defining the thickness of the intersection ($\gamma=$0.3 mm determined empirically). Ideally, $\mathcal{I}^{t}$ can be projected back onto the ultrasound image, forming a complete bone label. However, as also reported previously [20], mismatches (1 to 2 mm) often occur due to accumulated errors, including tracking error, calibration error (see Figure 2). Although these mismatches are typically below 2 mm, they are noticeable given the high resolution of ultrasound images, which can reach up to 0.05 mm per pixel.
@@
Bone contours that are roughly perpendicular to the ultrasound probe typically appear as high-intensity regions in ultrasound images due to the underlying ultrasound physics. Based on this principle, we propose an intensity-based optimization method designed to refine the location of CT-derived labels. The proposed optimization relies on the following assumptions: (i) the main error sources are tracking and calibration errors, as the only undetermined variables in Eq. (1) are the tracking matrices ($\mathsf{T}^{t+\Delta t}_{CT\leftarrow O}$ and $\mathsf{T}^{t+\Delta t}_{O\leftarrow US}$) and the calibration matrix ($\mathsf{T}_{US\leftarrow I}$); (ii) pixels with high intensity in the vicinity of the non-corrected labels mostly correspond to the true bone contour. Beneath the bone shadow, there are almost uniformly dark regions, as bones possess the highest acoustic impedance compared to the surrounding soft tissues, causing most acoustic waves to be reflected back; (iii) the CT-ultrasound alignment error does not exceed an mean distance of 2 mm, given that the optical tracking system offers sub-millimeter accuracy [^5] https://atracsys.com/fusiontrack-500/ and the calibration matrix achieves a 3D localization error of 1.07 mm. Furthermore, a random image sampled from each ultrasound sweep to measure mismatch yielded a mean error of $1.13\pm 0.606$ mm.

### v2

@@ Abstract
Background: Ultrasound-based bone surface segmentation is crucial in computer-assisted orthopedic surgery. However, ultrasound images have limitations, including a low signal-to-noise ratio, acoustic shadowing, and speckle noise, which make interpretation difficult. Existing deep learning models for bone segmentation rely primarily on costly manual labeling by experts, limiting dataset size and model generalizability. Additionally, the complexity of ultrasound physics and acoustic shadow makes the images difficult for humans to interpret, leading to incomplete labels in low-intensity and anechoic regions and limiting model performance. To advance the state of the art in ultrasound bone segmentation and establish effective model benchmarks, larger and higher-quality datasets are needed.
Methods: We propose a methodology for collecting ex-vivo ultrasound datasets with automatically generated bone labels, including anechoic regions. The proposed labels are derived by accurately superimposing tracked bone Computed Tomography (CT) models onto the tracked ultrasound images. These initial labels are refined to account for ultrasound physics. To clinically evaluate the proposed method, an expert physician from our university hospital specialized on orthopedic sonography assessed the quality of the generated bone labels. A neural network for bone segmentation is trained on the collected dataset and its predictions are compared to expert manual labels, evaluating accuracy, completeness, and F1-score.
Results: We collected the largest known dataset of 100k ultrasound images of human lower limbs with bone labels, called UltraBones100k. A Wilcoxon signed-rank test with Bonferroni correction confirmed that the bone alignment after our optimization pipeline significantly improved the quality of bone labeling ($p<0.001$). The model trained on UltraBones100k consistently outperforms manual labeling in all metrics, particularly in low-intensity regions (320% improvement in completeness at a distance threshold of 0.5 mm)
@@
Ultrasound bone segmentation , Bone surface segmentation , Bone surface reconstruction , Ultrasound image analysis , Computer-assisted orthopedic surgery
journal: Computers in Biology and Medicine
[…]
[label1]organization=Research in Orthopedic Computer Science, Balgrist University Hospital, University of Zurich,addressline=Lengghalde 5, city=Zurich, postcode=8008, state=Zurich, country=Switzerland
[…]
[label2]organization=Department of Orthopaedics, Balgrist University Hospital, University of Zurich,addressline=Forchstrasse 340, city=Zurich, postcode=8008, state=Zurich, country=Switzerland
[…]
[label3]organization=Computer Vision and Geometry Group, ETH Zurich,addressline=Ramistrasse 101, city=Zurich, postcode=8092, state=Zurich, country=Switzerland
@@ 1 Introduction
Ultrasound is a well-established imaging technology widely used in the clinical diagnosis of musculoskeletal pathologies, including bone and soft tissue injuries [1]. It also facilitates dynamic assessments of joint and tendon movements, stability[2], and the detection of fracture healing, structural abnormalities, infections, ligament injuries, nerve compression, and mechanical impingement [3]. As evidenced by a substantial body of literature [4, 5, 6, 7, 8], the automated segmentation and 3D reconstruction of bone surfaces from ultrasound images are gaining increasing significance as a powerful tool for interventional Computer-Assisted Orthopedic Surgery (CAOS) systems. Compared to ionizing imaging modalities like fluoroscopy, ultrasound offers advantages in improving guidance and navigation during interventional and surgical procedures[4], including trauma treatment[6], needle placement[5], and robotic spine surgery [9].
@@
Accurate 3D modeling of bone geometry requires precise 2D identification and segmentation of bone anatomies in ultrasound images, which relies heavily on high-quality labels produced by expert professionals[10, 11, 12, 13]. Due to the highest acoustic impedance compared to surrounding soft tissues, bones typically appear as high-intensity regions when the ultrasound probe is perpendicular to the bone surface. However, when the probe is positioned at a larger oblique angle to the bone surface, bones can appear as low-intensity regions as most portion of acoustic energy is reflected into surrounding soft tissues instead of back to the probe. These low-intensity regions can only be inferred by surgeons through anatomical knowledge but cannot be accurately labeled. The same applies to fully anechoic regions formed by the complete absorption of the ultrasound signal in the overlying bone layers.
@@
To address these challenges, various medical imaging methods based on handcrafted features have been developed to enhance the interpretation of ultrasound images [14, 15]. Recently, deep learning-based bone segmentation methods have shown promising improvements in robustness, accuracy, and processing speed [16, 17, 7, 18, 8, 19, 20]. However, the performance of such models relies on the quality of the manually labeled training dataset. Manual bone labeling is time-consuming, and requires specialized knowledge of ultrasound imaging and bone anatomy, which cannot be produced on a large scale if relying solely on human experts. The resulting datasets are often small, ranging from dozens to a few thousands labeled slices. This limits the performance and generalizability of the trained models. Additionally, most existing methods train and validate their models on non-public datasets and use non-standardized evaluation metrics, making it difficult to compare and benchmark the performance of different models across studies [21, 22]. The absence of large-scale, accurately labeled public datasets has been a critical obstacle for the community in developing more advanced models.
@@
In this work, we propose a methodology that enables the collection of large-scale ultrasound datasets with automatically generated bone labels. Instead of relying on manual labeling, our approach utilizes optical tracking to align a 3D bone model, extracted from a preoperative CT scan, with tracked ultrasound images. The initial alignment is further refined using our proposed intensity-based optimization strategy for each frame, rather than relying on traditional global optimization [23]. This approach enhances the precision and completeness of bone labels in ultrasound images. With this methodology, we collected a dataset of 100k annotated ultrasound images with bone labels from 14 human cadaveric lower legs. A deep learning model for bone segmentation is trained on the resulting UltraBones100k dataset, which significantly outperforms manual segmentation performed by an expert surgeon.
To ensure reproducibility of the results, the dataset and pretrained model weights, have been made publicly available [^1] Link will be available upon acceptance..
@@ 2 Related Work
Ultrasound Bone Segmentation. Numerous intensity-based bone segmentation methods have been developed [21, 22]. However, intensity-based methods face challenges in terms of robustness and accuracy. To mitigate these issues, frequency-based methods have been introduced, including phase symmetry and structured phase symmetry techniques [15, 14], which aim to enhance bone boundaries. With the advent of deep learning, neural network models such as CNNs, U-nets, and GANs have been employed for bone segmentation in ultrasound images, targeting various anatomical regions such as the spine [16, 17, 7], pelvis [18, 8], femur [19, 8, 20], tibia [8, 20, 19], fibula [19], and radius [20, 19]. These deep learning methods have demonstrated superior performance over traditional approaches in terms of accuracy, robustness, and runtime. For example, average bone surface distance errors have been reported to range from 0.2 mm to 1.7 mm [20, 24, 7, 18]. Based on accurate 2D bone segmentation results, the reconstructed bone surfaces achieve sub-millimetric one-sided CD [11]. However, these studies rely on manually annotated bone labels from expert surgeons. This process is time-consuming, costly, and requires specialized knowledge of ultrasound imaging and bone anatomy. In addition, different surgeons may also provide different labels for the same ultrasound image, leading to inconsistencies. As a result, the quality and the size of training datasets is limited, typically ranging from dozens to a few thousands of labeled images [22]. These constraints reduce model performance and generalizability. In addition, the absence of standardized evaluation metrics, with over 18 different metrics in use [21], further hinders effective model benchmarking and comparison.
Automatic labeling. Optical tracking markers have been applied in the medical imaging community for collecting datasets with automatically generated labels, such as tracked spine CT models [25] and surgical instruments [26]. So far, one previous work has attempted to automatically generate bone labels in ultrasound images using a tracked bone CT model [12]. However, their approach led to significant alignment errors due to error accumulation, ultimately requiring a fallback to manual labeling. Traditional image-based multi-modal registration methods, such as CT-US registration in Imfusion Suite  [^2] https://www.imfusion.com/, compute a transformation matrix to globally register one modality to another [23]. These methods are highly dependent on initialization and global intensity distribution, which can lead to local minima. As we will demonstrate, these methods do not provide sufficient accuracy for automatic labeling. In contrast, our proposed method optimizes the alignment for each frame locally, resulting in more accurate bone labeling.
@@ 3.1 Dataset Collection and Automatic Bone labeling
Acquisition Setup. Our dataset acquisition setup using human cadaveric specimens is illustrated in Figure 1. Ethics approval has been granted for our study by the Zurich cantonal ethics committee under BASEC Nr. 2023-01681. Prior to data acquisition, K-wires with a diameter of 2.5 mm and a length of 150 mm (DePuy Synthes, USA), along with custom-made 3D-printed marker mounts equipped with infrared-sensitive fiducial markers, were mounted to each specimen. The K-wires were used to stabilize the bones of interest, including the lower leg and foot, to prevent bone movement during scanning. Optical markers were tracked by an optical tracking system (FusionTrack 500, Atracsys, Switzerland). A CT scan was performed on each specimen with an image resolution of 512 × 512 pixels, an in-plane resolution of 0.839 mm × 0.839 mm, and a slice thickness of 0.6 mm (NAEOTOM Alpha, Siemens, Germany). Bone segmentation was generated using the region growing and thresholding functionalities in Mimics (Materialise, Leuven, Belgium). For data acquisition, each specimen was scanned nine times using a tracked ultrasound device (AixplorerUltimate, SuperSonic Imagine, France) equipped with a linear probe that has an imaging frequency range of 4.0 to 15.0 MHz. The scanning procedure consisted of three forward and backward passes along the longitudinal axis of the tibia, three along the fibula, and three covering the foot (including the bones of the ankle joint such as the calcaneus and talus), aiming to maximize bone surface coverage. Ultrasound images, along with the tracking data of the ultrasound probe and the specimen, were recorded with corresponding timestamps.
Figure 2: Diagram explaining the transformation chain across all components. Solid black arrows represent known transformations, which are directly measured by the optical tracking system. The dashed-line arrow represents the unknown transformation from the image plane to the ultrasound probe, which will be determined through calibration.
@@
System Calibration. To determine the coordinates of the tracking fiducials on the specimen marker within the CT space, virtual spheres are created to match the size of the physical fiducials in MeshLab[27]. These virtual spheres are initially manually aligned with the corresponding fiducials in the segmented CT model, followed by refinement using the point-to-point Iterative Closest Point algorithm [28]. The resulting registration transformation matrix allows for the accurate calculation of relative poses between the ultrasound probe and the bone CT model. However, converting each pixel of an ultrasound image into its corresponding point in the CT space requires accurate spatial calibration and accurate tracking of the ultrasound probe. In addition, temporal calibration is necessary to synchronize the ultrasound images with the tracking data. Following the convention in existing works [29, 30], we define the origin of the ultrasound image at the top-left corner. The transformation chain across all components is illustrated in Figure 2. The pixels of the ultrasound image at timestamp $t$ are defined in homogeneous coordinates as $\mathcal{P}^{t}=\left\{\begin{bmatrix}u,v,0,1\end{bmatrix}^{T}\mid u\in[0,W),v\in[0,H)\right\}$, where $W$ and $H$ represent the image width and height, respectively. The corresponding point cloud in CT space, $\mathcal{U}^{t}=\left\{\begin{bmatrix}x_{j},y_{j},z_{j},1\end{bmatrix}^{T}\mid j\in[0,H\times W)\right\}$, is computed as
$\mathcal{U}^{t}=\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{OT}}\cdot\mathsf{T}^{t+\Delta t}_{\text{OT}\leftarrow\text{US}}\cdot\mathsf{T}_{\text{US}\leftarrow\text{IP}}\cdot\mathsf{T}_{\text{S}}\cdot\mathcal{P}^{t}=\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{US}}\cdot\mathsf{T}_{\text{US}\leftarrow\text{IP}}\cdot\mathsf{T}_{\text{S}}\cdot\mathcal{P}^{t}$ (1)
with:
  • 1.
    $\mathsf{T}_{\text{S}}=\begin{bmatrix}s_{x}&0&0&0\\ 0&s_{y}&0&0\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}$, where $s_{x}$ is the width of one pixel in mm, and $s_{y}$ is the height of one pixel in mm. $s_{x}$ and $s_{y}$ are known from the ultrasound device.
  • 2.
    $\mathsf{T}_{\text{US}\leftarrow\text{IP}}=\begin{bmatrix}r_{11}&r_{12}&r_{13}&d_{x}\\ r_{21}&r_{22}&r_{23}&d_{y}\\ r_{31}&r_{32}&r_{33}&d_{z}\\ 0&0&0&1\end{bmatrix}$, belonging to the the 3D Special Euclidean group $\text{SE}(3)$, is the transformation matrix that maps 3D points from the coordinate system of the image plane (origin: top-left corner) to the coordinate system of the marker attached to the ultrasound probe. The determination of this matrix refers to the process of spatial ultrasound calibration [30, 29]. Note that a transformation matrix $\mathsf{T}(\bm{\theta},\mathbf{d})$ can be defined using the Euler rotation angles $\bm{\theta}$ and the translation vector $\mathbf{d}$, which together specify the relationship between two coordinate systems.
  • 3.
    $\mathsf{T}^{t+\Delta t}_{\text{OT}\leftarrow\text{US}}\in\text{SE}(3)$ is the transformation matrix at timestamp $t+\Delta t$ that transforms 3D points from the coordinate system of the ultrasound probe marker to the coordinate system of the optical tracking system. As the probe moves during dataset collection, $\mathsf{T}_{O\leftarrow US}$ changes over time instead of a constant matrix. At each timestamp $t$, two types of data were collected: ultrasound image data ($\mathcal{P}^{t}$) and tracking data from multiple markers ($\mathsf{T}^{t}$). Due to differences in sensor characteristics, such as frame rates and data acquisition timings, $\mathcal{P}^{t}$ and $\mathsf{T}^{t}$ are not inherently synchronized, resulting in a temporal offset denoted as $\Delta t$. Determining $\Delta t$ is thus the process of temporal calibration. In our setup, the tracking data has a higher temporal resolution than the ultrasound data ($\Delta t<0$), as forming a B-mode ultrasound image requires additional time to process the original acoustic signals.
  • 4.
    $\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{OT}}\in\text{SE}(3)$ is the transformation matrix at timestamp $t+\Delta t$ from the coordinate system of the optical tracking system to the coordinate system of the CT scan, which can be determined using the marker attached to the cadaveric specimen.
  • 5.
    For simplicity, we denote $\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{US}}(\bm{\theta},\mathbf{d})=\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{OT}}(\bm{\theta}_{1},\mathbf{d}_{1})\cdot\mathsf{T}^{t+\Delta t}_{\text{OT}\leftarrow\text{US}}(\bm{\theta}_{2},\mathbf{d}_{2})$ as the direct transformation from the ultrasound marker to the CT space.
@@
Unlike previous studies that treated spatial and temporal calibration independently [30, 29], we performed both calibrations simultaneously using a Z-phantom [30]. The temporal offset $\Delta t$ was sampled over the range $[-300,0)$ ms with a step size of 0.1 ms. For each $\Delta t$, $\mathsf{T}_{\text{US}\leftarrow\text{IP}}$ was computed, and the corresponding 3D localization recorded. The optimal parameters were identified by minimizing the 3D localization error. Our method achieved a 3D localization error of 1.07 mm, outperforming previously reported errors ($\approx 1.5$ mm) [30, 29]. With precise calibration results, 2D pixel coordinates can be accurately mapped to 3D points in the CT space. For simplicity, we denote the pixel corresponding to the point $\mathbf{u}_{i}^{t}\in\mathcal{U}^{t}$ as $\mathbf{p}_{i}^{t}\in\mathcal{P}^{t}$, and define $f(\mathbf{u}_{i}^{t})$ as the intensity value of $\mathbf{p}_{i}^{t}$ in the original ultrasound image.
Initial Bone Labeling Points are sampled uniformly over the bone CT model to generate the CT point cloud $\mathcal{C}$. Both $\mathcal{C}$ and $\mathcal{U}^{t}$ (with original intensity values) are rendered in Figure 3. The CT-ultrasound bone surface intersection is defined as $\mathcal{I}^{t}=\mathcal{C}\cap\mathcal{U}^{t}=\left\{\mathbf{u}\in\mathcal{U}^{t}\mid\min_{\mathbf{c}\in\mathcal{C}}\lVert\mathbf{u}-\mathbf{c}\rVert\leq\gamma\right\}$ where $\gamma$ is the parameter defining the thickness of the intersection ($\gamma=$0.3 mm determined empirically). Ideally, $\mathcal{I}^{t}$ can be projected back onto the ultrasound image, forming a complete bone label. However, as also reported previously [12], mismatches (1 to 2 mm) often occur due to accumulated errors, including tracking error and calibration error (see Figure 3). Although these mismatches are typically below 2 mm, they are noticeable given the high resolution of ultrasound images, which can reach up to 0.05 mm per pixel.
@@
Optimization Step for Refined Alignment. Bone contours that are roughly perpendicular to the ultrasound probe typically appear as high-intensity regions in ultrasound images due to the underlying ultrasound physics. Based on this principle, we propose an intensity-based optimization method designed to refine the location of CT-derived labels. The proposed optimization relies on the following assumptions: (i) the main error sources are tracking and calibration errors, as the only undetermined variables in Eq. (1) are the tracking matrices ($\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{OT}}$ and $\mathsf{T}^{t+\Delta t}_{\text{OT}\leftarrow\text{US}}$) and the calibration matrix ($\mathsf{T}_{\text{US}\leftarrow\text{IP}}$); (ii) pixels with high intensity in the vicinity of the non-corrected labels mostly correspond to the true bone contour. Beneath the bone shadow, there are almost uniformly dark regions, as bones possess the highest acoustic impedance compared to the surrounding soft tissues, causing most acoustic waves to be reflected back; (iii) the CT-ultrasound alignment error does not exceed an mean distance of 2 mm, given that the optical tracking system offers sub-millimeter accuracy [^3] https://atracsys.com/fusiontrack-500/ and the calibration matrix achieves a 3D localization error of 1.07 mm. Furthermore, a random image sampled from each ultrasound sweep to measure mismatch yielded a mean error of $1.13\pm 0.606$ mm.